OpenStudy (anonymous):

Why everyone got to hate? D: help please D:::: Find 3 pairs of algebraic expressions, each pair having a product of a^5.

OpenStudy (jamesj):

So you want a pair of expressions, call them X1 and X2, such that \[ X_1 \cdot X_2 = a^5 \] ?

OpenStudy (anonymous):

\[a^2\times a^3\] could be one of them. bet you can find more pairs of numbers that add to 5

OpenStudy (jamesj):

How about to begin \[ X_1 = a^2, \ X_2 = a^3 \]

OpenStudy (jamesj):

great minds ...

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Thats a pair? I thought it was something else...oh well, thank you

OpenStudy (jamesj):

\[ X_1 = \frac{a}{b^{1024}}, \ C_2 = a^4b^{1024} \]

OpenStudy (anonymous):

thank you . :D even though i cant read it... its in code form, but thanks

OpenStudy (jamesj):

..and that should be X_2. But you get the idea.